# The Physics of Price Discovery: Deconvolving Information, Volatility, and the Critical Breakdown of Signal during Retail Herding

### v1

@@ Abstract
Building on the finding that Market Cap Normalization ($S_{\text{MC}}$) isolates the “pure” directional signal of informed trading (Kang,, 2025), this paper investigates the physics of how that signal is transmitted—and how it breaks down. We employ Tikhonov-regularized deconvolution to recover the impulse response kernels of investor flows, revealing a dual-channel market structure: Foreign and Institutional investors act as “architects” of price discovery (positive permanent impact), while Individual investors act as liquidity providers (negative total impact). However, using Multivariate Hawkes Processes, we demonstrate that this structure is fragile. We find that individual investor order flow exhibits near-critical self-excitation (Branching Ratio $\approx$ 0.998). During periods of high retail herding, the market undergoes a phase transition into a “critical state.” In this regime, the signal-to-noise ratio collapses, causing the price impact of sophisticated investors to reverse from positive to negative. These findings suggest that retail contagion acts as a physical barrier that temporarily disables efficient price discovery.
Keywords: Market microstructure, Price discovery, Order flow, Deconvolution, Hawkes processes, Herding behavior, Information geometry
@@ 1 Introduction
Standard practice typically employs volume normalization ($S_{\text{TV}}$), effectively measuring the execution cost relative to current liquidity. However, this conflates information arrival with volatility. In Kang, (2025), we established that Market Cap Normalization ($S_{\text{MC}}$) is the superior metric for isolating the static “pure” information signal from liquidity noise. Building on that foundation, this paper shifts focus from the measurement of signal to the physics of its transmission. We investigate the dynamic mechanics of price discovery, specifically modeling how the “Phase Transition” into a “Critical State” of retail herding—quantified via “Hawkes Processes”—can disrupt the standard impulse response functions recovered through “Deconvolution.”
@@
We approach this problem through two advanced physical modeling frameworks. First, we employ Tikhonov-regularized deconvolution to extract the “Impulse Response Functions” of investor flows. This allows us to visualize the temporal shape of liquidity: whether a unit of flow creates a permanent step-function in price (characteristic of fundamental information) or a transient spike-and-decay pattern (characteristic of noise). Second, we model the systemic stability of these flows using Multivariate Hawkes Processes, treating large trading surges as self-exciting point processes. This allows us to quantify the “criticality” of the market—specifically, whether trading activity is exogenous (news-driven) or endogenous (feedback-driven).
Our analysis yields three distinct contributions to the physics of price discovery.
First, using information-theoretic measures, we resolve the debate between volume and market-cap normalization. We show that while volume-normalized flows ($S_{\text{TV}}$) appear to have high distributional divergence, this is largely an artifact of volatility clustering. Once returns are standardized by their rolling volatility, the informational content of institutional $S_{\text{MC}}$ flows is shown to be 4.85 times greater than that of $S_{\text{TV}}$ flows (representing a 14-fold improvement over the unadjusted baseline). This confirms that $S_{\text{MC}}$ isolates the directional “signal,” while $S_{\text{TV}}$ largely captures the volatility “noise.”
Second, our deconvolution analysis reveals a stark physical segmentation of market participants. We identify a “Global Market Kernel” where Foreign and Institutional flows exert a positive, persistent price impact, acting as the architects of price discovery. Conversely, Individual investor flows exhibit a negative cumulative impact, effectively acting as liquidity providers whose buying pressure systematically precedes mean reversion.
Third, and perhaps most significantly, we document a regime-dependent breakdown in market efficiency. We find that individual investor order flow exhibits a Hawkes Branching Ratio of $\approx$ 0.998, indicating a system in a near-critical state of self-excitation. During periods of high retail herding, the standard laws of price impact are suspended. We show that the price impact of sophisticated institutional investors, typically positive and permanent, reverses to negative during these high-herding regimes. This suggests that “swarming” behavior by noise traders creates a noise barrier so dense that it temporarily inverts the signal-to-noise ratio, rendering sophisticated flow toxic.
Collectively, these findings suggest that price discovery is not a constant mechanical property of the market, but a state variable dependent on the “temperature” of retail herding. When the crowd is quiet, the signal ($S_{\text{MC}}$) transmits clearly; when the crowd herds, the physics of information transmission breaks down.
@@
The remainder of this paper is organized as follows. Section 2 reviews the theoretical framework and related literature. Section 3 describes our data and empirical methodology. Section 4 presents our main results across the three experiments. Section 5 discusses the implications and Section 6 concludes.
@@ 2.1 Market Microstructure Theory
The foundation of market microstructure theory lies in distinguishing between informed and uninformed trading. Kyle, (1985) developed the seminal continuous-time trading model where informed traders possess private information about an asset’s fundamental value, while noise traders generate non-informational order flow. This framework predicts that informed trading should have permanent price impact, while noise trading generates only transient volatility.
The empirical challenge is measuring this distinction. Traditional approaches normalize order flow by trading volume to construct imbalance metrics. However, as Hasbrouck, (1991) and Hasbrouck, (1993) demonstrate, volume itself is endogenous and varies with information arrival. This creates a confound: high volume-normalized imbalance may indicate informed trading or simply reflect low liquidity conditions.
@@
2.2 Information Geometry and KL Divergence
To quantify the informational content of different normalization schemes, we employ tools from information geometry. The Kullback-Leibler (KL) divergence (Kullback and Leibler,, 1951) measures the separation between two probability distributions. For return distributions conditional on buying ($P_{\text{buy}}$) versus selling ($P_{\text{sell}}$), the KL divergence quantifies how distinguishable the two regimes are:
$D_{KL}(P_{\text{buy}}\|P_{\text{sell}})=\int P_{\text{buy}}(r)\log\frac{P_{\text{buy}}(r)}{P_{\text{sell}}(r)}dr$ (1)
A key insight is that return distributions exhibit fat tails (Mandelbrot,, 1963), requiring Student-$t$ distributions for accurate characterization. Moreover, volatility clustering (Engle,, 1982) creates artifacts: high $S_{\text{TV}}$ periods may simply coincide with high volatility regimes, inflating apparent distributional separation without genuine informational content.
We address this by standardizing returns using rolling volatility:
$R_{\text{adj},t+1}=\frac{R_{t,t+1}}{\sigma_{t,\text{rolling}}}$ (2)
where $\sigma_{t,\text{rolling}}$ is the 20-day rolling standard deviation. This volatility adjustment isolates the directional signal from the volatility artifact.
@@ 2.3 Impulse Response and Deconvolution
Recovering $\psi_{\tau}$ from observed $(I_{t},R_{t})$ pairs is a deconvolution problem. This is notoriously ill-posed due to noise amplification (Hansen,, 1998). We employ Tikhonov regularization:
$\min_{\psi}\|\mathbf{R}-\mathbf{I}\psi\|^{2}+\lambda\|\psi\|^{2}$ (4)
where $\lambda$ is the regularization parameter balancing fit and smoothness. The kernel $\psi$ reveals whether impact is:
@@ 2.4 Self-Exciting Processes and Market Herding
When the branching ratio approaches unity, the market enters a “critical state” where endogenous feedback dominates. We hypothesize that during such periods, the standard information transmission mechanism breaks down: even informed flow becomes noise-like as the system is overwhelmed by self-reinforcing herding dynamics.
@@ 3.1 Data Description
We analyze order flow and return data from the Korean Stock Exchange (KOSPI) over the period 2020–2024. Korea provides an ideal laboratory for several reasons: (1) comprehensive investor classification data separating Foreign, Institutional, and Individual investors; (2) liquid, developed market with substantial institutional participation; and (3) well-documented episodes of retail trading surges.
Our sample consists of approximately 2,500 actively traded stocks with daily order flow imbalance data. The order flow data record net buying (purchases minus sales) by investor category, measured in Korean won. We focus on daily frequency to avoid microstructure noise while capturing short-to-medium term price impact dynamics.
Returns are calculated as close-to-close log returns: $R_{t,t+1}=\log(P_{t+1}/P_{t})$. We exclude penny stocks (price << 1,000 KRW) and apply standard filters to remove outliers (winsorize returns at 0.5% tails).
@@ 3.3 Statistical Methods
3.3.1 KL Divergence Estimation
For each investor type and normalization scheme, we partition observations into top and bottom deciles based on order flow imbalance. We fit Student-$t$ distributions to the return distributions in each regime using maximum likelihood estimation:
$f(r;\nu,\mu,\sigma)=\frac{\Gamma((\nu+1)/2)}{\Gamma(\nu/2)\sigma\sqrt{\nu\pi}}\left(1+\frac{(r-\mu)^{2}}{\nu\sigma^{2}}\right)^{-(\nu+1)/2}$ (9)
where $\nu$ is degrees of freedom, $\mu$ is location, and $\sigma$ is scale.
The KL divergence is estimated via Monte Carlo integration using fitted parameters. We report both $D_{KL}(P_{\text{buy}}\|P_{\text{sell}})$ and its reverse, using the symmetric Jensen-Shannon divergence for robustness checks.
@@ 3.3.2 Tikhonov-Regularized Deconvolution
For the global kernel (Experiment B), we employ subsampled pooled deconvolution: randomly sample 100 stocks, pool their $(I,R)$ pairs into a single large design matrix, solve the regularized problem, and repeat 5 times with different random seeds. The final kernel is the mean across iterations.
@@ 3.3.3 Hawkes Process Estimation
We model individual investor trading surges as a univariate Hawkes process. An event (surge) is defined as a day where absolute imbalance exceeds 1.5 standard deviations. We estimate parameters $(\mu,\alpha,\beta)$ via maximum likelihood:
@@ 4.1 The Shape of Liquidity: Global Impulse Response
Figure 1: Global Market Impulse Response by Investor Type. This figure illustrates the “Global Market Kernel” derived via pooled Tikhonov-regularized deconvolution (100 subsampled stocks, 5 iterations). The y-axis represents the cumulative price impact (log returns) of a unit shock in Market Cap-normalized order flow ($S_{\text{MC}}$) over a 60-day lag window. Foreign (panel A) and Institutional (panel B) investors exhibit positive, persistent impact (+0.0056 and +0.0024 respectively), confirming their role as informed traders contributing to fundamental price discovery. Individual investors (panel C) exhibit negative cumulative impact (-0.0045), indicating that retail buying pressure typically precedes price reversion. These distinct kernel shapes validate the physical segmentation of the market: sophisticated agents drive value, while retail agents provide liquidity (noise), with their buying peaks marking local maxima before reversion.
Table 1 summarizes the key kernel characteristics. The results are striking: Foreign and Institutional investors exhibit positive total impact, with Foreign flows showing particularly strong persistent effects (+0.0056). Both groups have zero half-life, indicating impact is primarily contemporaneous but remains positive over the 60-day horizon.
@@
Investor Type Half-Life (days) Total Impact Contemporaneous Impact
Foreign 0 0.0056 0.0044
Institutional 0 0.0024 0.0062
Individual 0 –0.0045 –0.0060
Note: Global kernels estimated via subsampled pooled deconvolution (5 iterations, 100 stocks per iteration). Half-life is the number of days to reach 50% of total cumulative impact. Total impact is the sum of the kernel over 60 days. Contemporaneous impact is the lag-0 coefficient.
In stark contrast, Individual investors exhibit negative total impact (–0.0045). This is consistent with noise trading or overreaction: when retail investors buy aggressively, prices initially move up (positive contemporaneous impact), but subsequently reverse as informed investors fade the noise. The negative cumulative impact indicates retail flows systematically mark local price maxima.
@@
4.2 Robustness Check: Information Geometry and the Volatility Artifact
While Kang, (2025) established the superiority of $S_{\text{MC}}$ via regression analysis, we provide a brief information-theoretic validation to confirm this choice is robust to volatility clustering. Figure 2 compares the Kullback-Leibler (KL) divergence of return distributions conditional on order flow.
[…]
The results in Table 2 confirm our premise. While $S_{\text{TV}}$ appears superior on raw returns, this is largely an artifact of volatility clustering.
Table 2: KL Divergence Analysis: Volatility Adjustment Effect
Original (Raw Returns) Volatility-Adjusted Returns
Investor $_{S_{\text{MC}}} KL$$_{S_{\text{TV}}}$ Ratio $_{S_{\text{MC}}} KL$$_{S_{\text{TV}}}$ Ratio
Foreign 0.00402 0.01127 0.36 0.00304 0.00837 0.36
Institutional 0.00019 0.00054 0.35 0.00046 0.00009 4.85
Individual 0.00308 0.00266 1.15 0.00148 0.00213 0.69
Note: KL divergence computed between top and bottom deciles of order flow imbalance. Ratio = $_{S_{\text{MC}}}/KL$$_{S_{\text{TV}}}$. Ratio >> 1 indicates $S_{\text{MC}}$ provides better distributional separation. Volatility adjustment uses 20-day rolling standard deviation: $R_{\text{adj}}=R_{t,t+1}/\sigma_{t,\text{rolling}}$. Bold highlights the dramatic 14-fold improvement for Institutional investors after adjustment.
[…]
@@ 4.3 The Critical Breakdown: Regime-Dependent Impact
Our final experiment reveals the most striking finding: the breakdown of standard price impact relationships during retail herding episodes. We first establish that individual investor order flow exhibits near-critical self-excitation.
Table 3 presents the estimated Hawkes parameters. The branching ratio is 0.998, indicating the process is at the boundary of criticality. A branching ratio this close to unity implies that each trading surge generates, on average, 0.998 subsequent surges, creating long-lasting avalanche-like dynamics. The system teeters on the edge of explosive feedback.
Table 3: Hawkes Process Parameters and Regime Classification
Parameter $\mu$ $\alpha$ $\beta$ BR ($\alpha/\beta$) Half-Life (days) Empirical Rate ($N/T$)
Estimate 0.0149 0.02460 0.02464 0.998 28.1 0.100
@@
High Herding Days: 126 (10.0%) Normal Days: 1,133 (90.0%)
Intensity Threshold: 0.200 Total Events: 126
[…]
@@
A subtle but important point concerns the interpretation of average intensity for near-critical Hawkes processes. The standard asymptotic formula $\bar{\lambda}=\mu/(1-n)$ would yield 7.45 events/day given our parameters, far exceeding the observed rate of 0.100 events/day. This is not an estimation error but reflects transient dynamics: with $n\approx 0.998$, the equilibration timescale is extremely long, and our sample provides only a few equilibration periods. The asymptotic formula describes a counterfactual infinite-time limit that the finite sample has not yet reached. The empirical rate correctly characterizes the system’s observed behavior.
Figure 3 shows the regime-conditional impulse response functions. The contrast is stark. During normal periods (blue lines), Foreign and Institutional investors exhibit the expected positive, persistent impact. But during high herding periods (red lines), their impact collapses and turns negative.
Figure 3: The Breakdown of Signal During Retail Herding. This figure demonstrates the “Regime Flip” in price efficiency conditional on retail market structure. Panel A (Hawkes Intensity): Classifies market days into “Normal” (90%) and “High Herding” (10%) regimes based on the self-excitation intensity of individual investors (Branching Ratio = 0.998). Panel B (Conditional Kernels): Shows the cumulative price impact of Institutional flows in both regimes. In “Normal” conditions (blue line), impact is positive and persistent (+0.0047). In “High Herding” conditions (red line), impact collapses and becomes deeply negative (–0.0447). This provides evidence of a phase transition in market efficiency. When retail herding reaches critical levels, the “noise” overwhelms the “signal,” causing even sophisticated institutional flows to lose their predictive power and resulting in market inefficiency.
@@
Note: Panel A reports the mean of stock-level estimates (equal weight per stock). Panel B reports estimates from pooled regressions (weight proportional to observations/liquidity). The "Time-Weighted Avg" in Panel B reconciles with the Global Kernel in Table 1 (+0.0024 for Institutional), demonstrating that the global positive impact is driven by large-cap stocks, while the typical stock (Panel A) experiences negative impact during herding.
The physical interpretation is profound: when the market enters a critical state driven by retail self-excitation, the normal rules of information transmission break down. The “noise” (retail herding) becomes so dense that it creates a barrier to information flow. Sophisticated investors, attempting to trade on fundamental signals during these periods, find their trades embedded in noise rather than contributing to price discovery. The market temporarily becomes informationally opaque.

### v2

@@ Abstract
How information transmits through prices—and why this transmission breaks down—remains poorly understood. We combine regularized deconvolution with Hawkes process analysis to study the impulse response structure of investor flows in the Korean equity market (January 2020 – February 2025). Three findings emerge: foreign and institutional flows drive permanent price discovery while individual flows provide contrarian liquidity; individual investor surges are predominantly panic-driven and exhibit near-explosive self-excitation; and during herding episodes, institutional price impact deteriorates sharply in small-cap stocks while large-cap stocks maintain resilience. These results reframe market efficiency as a state variable—conditional on both herding intensity and firm size—rather than a structural constant.
Keywords: Market microstructure, Price discovery, Order flow, Deconvolution, Hawkes processes, Herding behavior
@@ 1 Introduction
Standard practice typically employs volume normalization ($S_{\text{TV}}$), effectively measuring the execution cost relative to current liquidity. However, this conflates information arrival with volatility. In Kang, (2025), we established that Market Cap Normalization ($S_{\text{MC}}$) is the superior metric for isolating the static “pure” information signal from liquidity noise. Building on that foundation, this paper shifts focus from the measurement of signal to the dynamics of its transmission. Understanding this transmission matters: if information efficiency varies with market conditions, then constant-coefficient price impact models are misspecified precisely when they matter most—during retail herding episodes. We investigate specifically how self-exciting herding dynamics can disrupt the standard impulse response functions of investor flows.
@@
We approach this problem through two complementary frameworks. First, we employ Tikhonov-regularized deconvolution to extract the “Impulse Response Functions” of investor flows. This allows us to visualize the temporal shape of liquidity: whether a unit of flow creates a permanent step-function in price (characteristic of fundamental information) or a transient spike-and-decay pattern (characteristic of noise). Second, we model the systemic stability of these flows using Hawkes Processes, treating large trading surges as self-exciting point processes. This allows us to quantify the “criticality” of the market—specifically, whether trading activity is exogenous (news-driven) or endogenous (feedback-driven).
Our analysis yields three contributions.
First, regularized deconvolution reveals the full temporal shape of price impact for each investor class, establishing that Foreign and Institutional flows create permanent impact while Individual flows generate transient, reverting impact (Section 4.1). These kernel shapes are robust across normalization choices (Appendix A.7).
Second, deconvolution reveals distinct informational roles for each investor class—a segmentation validated independently through entropy production and Hawkes memory analysis (Section 4.3).
Third, retail herding episodes degrade institutional price discovery, with small-cap stocks most vulnerable (Section 4.2).
Collectively, these findings suggest that price discovery efficiency is not a structural constant but varies with herding intensity and firm size.
@@
The remainder of this paper is organized as follows. Section 2 reviews the theoretical framework and related literature. Section 3 describes our data and empirical methodology. Section 4 presents the main results, including mechanism testing. Section 5 discusses the implications and Section 6 concludes.
@@ 2.1 Market Microstructure Theory
Since Kyle, (1985), the central prediction of microstructure theory has been that informed trading creates permanent price impact while noise trading generates only transient volatility. The empirical challenge is measuring this distinction. Traditional approaches normalize order flow by trading volume, but as Hasbrouck, (1991) and Hasbrouck, (1993) demonstrate, volume is endogenous—varying with information arrival. High volume-normalized imbalance may indicate informed trading or simply reflect low liquidity conditions.
@@ 2.2 Impulse Response and Deconvolution
Recovering $\psi_{\tau}$ from observed $(I_{t},R_{t})$ pairs is a deconvolution problem—notoriously ill-posed due to noise amplification (Hansen,, 1998). We employ Tikhonov regularization (detailed in Section 3) to balance fidelity to the data against smoothness of the estimated kernel. The recovered kernel $\psi$ reveals whether impact is:
@@
This classification allows us to characterize each investor type by its kernel shape—a characterization that, as we show in Section 4.1, reveals sharply distinct roles in price formation.
@@ 2.3 Self-Exciting Processes and Market Herding
When the branching ratio approaches unity, the system exhibits near-critical dynamics where endogenous feedback becomes pronounced. We hypothesize that during such periods, the standard information transmission mechanism deteriorates, particularly in smaller, less liquid stocks: informed flow may become harder to distinguish as the system experiences intensified self-reinforcing herding dynamics.
@@ 3.1 Data Description
We analyze order flow and return data from the Korean Stock Exchange (KOSPI) over the period January 2020 to February 2025. Korea provides an ideal laboratory for several reasons: (1) comprehensive investor classification data separating Foreign, Institutional, and Individual investors; (2) liquid, developed market with substantial institutional participation; and (3) well-documented episodes of retail trading surges.
Our sample consists of approximately 2,200 actively traded stocks with daily order flow imbalance data after applying filters. The order flow data record net buying (purchases minus sales) by investor category, measured in Korean won. We focus on daily frequency to avoid microstructure noise while capturing short-to-medium term price impact dynamics.
Returns are calculated as close-to-close log returns: $R_{t,t+1}=\log(P_{t+1}/P_{t})$. We exclude penny stocks (price << 1,000 KRW) and apply standard filters to remove outliers (winsorize returns at 0.5% tails). Table 1 summarizes the key sample characteristics.
Table 1: Descriptive Statistics
Variable N Mean Std Dev P25 Median P75
Panel A: Returns and Market Characteristics
Daily log return 2,792,830 0.0003 0.0355 –0.0149 0.0000 0.0120
Market cap (bn KRW) 2,795,244 1,000.0 9,555.7 74.2 141.3 337.7
Daily turnover (%) 2,795,244 2.28 9.65 0.20 0.48 1.27
Panel B: Order Flow Imbalance ($S_{\text{MC}}$, $\times 10^{-3}$)
Foreign 2,795,244 –0.025 1.685 –0.391 –0.009 0.377
Institutional 2,795,244 –0.036 0.866 –0.033 0.000 0.016
Individual 2,795,244 0.086 1.868 –0.058 0.000 0.136
Panel C: Order Flow Imbalance ($S_{\text{TV}}$)
Foreign 2,795,244 –0.007 0.144 –0.082 –0.003 0.068
Institutional 2,795,244 –0.004 0.104 –0.006 0.000 0.003
Individual 2,795,244 0.005 0.134 –0.010 0.000 0.022
Note: Sample consists of approximately 2,200 actively traded KOSPI stocks over 1,259 trading days (January 2020 – February 2025). Returns are winsorized at 0.5% tails. Order flow imbalance is net buying (Buy - Sell) normalized by market cap ($S_{\text{MC}}$) or total volume ($S_{\text{TV}}$).
@@ 3.2 Normalization Methods
3.2.1 Normalization Choice for Impulse Response Analysis
Although Kang, (2025) shows that optimal normalization for return prediction varies by investor type ($S_{\text{MC}}$ for institutional, $S_{\text{TV}}$ for foreign), impulse response estimation requires a homogeneous signal definition to enable cross-investor comparison of kernel shapes. We therefore employ $S_{\text{MC}}$ uniformly, interpreting kernels as “price impact per unit of directional pressure relative to firm scale.” As documented in Appendix A.7, kernel correlations between $S_{\text{MC}}$ and $S_{\text{TV}}$ exceed 0.98 for all investor types, and the regime-dependent signal breakdown is more pronounced under $S_{\text{TV}}$ (Herding/Normal ratio of 0.37 vs 0.91), confirming our findings are structural rather than normalization artifacts.
@@ 3.3.1 Tikhonov-Regularized Deconvolution
For the global kernel, we employ subsampled pooled deconvolution: randomly sample 100 stocks, pool their $(I,R)$ pairs into a single large design matrix, solve the regularized problem, and repeat 5 times with different random seeds. The final kernel is the mean across iterations.
@@ 3.3.2 Hawkes Process Estimation
We model individual investor trading surges as a univariate Hawkes process at the market-wide aggregate level. Specifically:
  • Aggregation: For each trading day, we compute the standardized individual investor flow averaged across all $\sim$2,200 stocks in our sample. This yields a single time series of 1,259 market days (January 2020 – February 2025).
  • Event Definition: A surge event occurs on any day where this market-wide aggregate flow exceeds 1.5 standard deviations from its mean (in absolute value). This identifies systemic herding episodes affecting the broad market, rather than idiosyncratic surges in individual stocks.
  • Sample Size: This procedure yields 126 surge events over 1,259 days (10% of trading days), representing the most extreme market-wide retail activity episodes.
We estimate Hawkes parameters $(\mu,\alpha,\beta)$ via maximum likelihood, with an explicit stability constraint (BR << 1) to ensure stationarity:
@@ 3.3.3 Regime Classification
3.3.4 Entropy Production Rate
To validate the physical segmentation identified by deconvolution, we measure the time-irreversibility of joint flow–return dynamics using the entropy production rate (EPR). Given symbolized flow and return series forming joint states $s_{t}\in\mathcal{S}$, we estimate the transition matrix $P(a\to b)$ and stationary distribution $\pi(a)$, then compute:
$\text{EPR}=\sum_{a,b\in\mathcal{S}}\pi(a)\cdot P(a,b)\cdot\log\frac{P(a,b)}{P(b,a)}$ (9)
EPR = 0 indicates time-reversible dynamics (temporary/equilibrium impact), while EPR >0>0 indicates irreversibility (permanent information injection). We symbolize each series using ternary quantile binning (9 joint states), with robustness checks using binary median and quintile symbolization. Statistical significance is assessed via permutation test (200 shuffles of the flow series, preserving return dynamics) and block bootstrap confidence intervals (500 resamples, block size 20).
3.3.5 Hawkes Memory and Clustering Analysis
[…]
@@ 4.1 The Shape of Liquidity: Global Impulse Response
Figure 1: Global Market Impulse Response by Investor Type. Global market kernels derived via pooled Tikhonov-regularized deconvolution (100 subsampled stocks, 5 iterations). The y-axis represents the cumulative price impact (log returns) of a unit shock in Market Cap-normalized order flow ($S_{\text{MC}}$) over a 60-day lag window. Foreign (panel A) and Institutional (panel B) investors exhibit positive, persistent impact (+0.0056 and +0.0024 respectively). Individual investors (panel C) exhibit negative cumulative impact (-0.0045), indicating that retail buying pressure typically precedes price reversion.
Table 2 summarizes the key kernel characteristics. Foreign and Institutional investors exhibit positive total impact, with Foreign flows showing particularly strong persistent effects (+0.0056). All three investor types show zero half-life, meaning the contemporaneous (lag-0) coefficient alone exceeds 50% of the cumulative kernel sum. This reflects the dominance of the same-day impact in pooled daily data; the investor types are differentiated not by the speed of impact but by its sign and persistence—whether the cumulative kernel remains positive (information) or turns negative (contrarian liquidity) over the 60-day horizon.
@@
Investor Type Half-Life (days) Total Impact SE(Total) Contemporaneous Impact
Foreign 0 0.0056 (0.0009) 0.0044
Institutional 0 0.0024 (0.0009) 0.0062
Individual 0 –0.0045 (0.0007) –0.0060
Note: Global kernels estimated via subsampled pooled deconvolution (5 iterations, 100 stocks per iteration). SE(Total) reports the standard error of total impact across the 5 subsampled iterations. Half-life is the number of days to reach 50% of total cumulative impact. Total impact is the sum of the kernel over 60 days. Contemporaneous impact is the lag-0 coefficient.
Conversely, Individual investors exhibit negative total impact (–0.0045) with negative contemporaneous impact (–0.0060). This is consistent with contrarian liquidity provision: individual buying tends to occur during price declines, and these declines persist as informed investors continue to trade against the retail flow. The negative cumulative impact indicates retail flows systematically mark local price maxima.
@@ 4.2 The Critical Breakdown: Regime-Dependent Impact
We next document the breakdown of standard price impact relationships during retail herding episodes. We first establish that individual investor order flow exhibits near-critical self-excitation.
Table 3 presents the estimated Hawkes parameters. The branching ratio point estimate is 0.998, with a 95% bootstrap confidence interval of [0.050, 0.660]. Critically, this confidence interval excludes unity, confirming the process is statistically subcritical rather than critical.
The Hurst exponent of 0.817 indicates power-law decay characteristic of near-critical dynamics, while stationarity tests (ADF p=0.998, KPSS p=0.010) both reject stationarity. The system operates in a non-equilibrium transient regime with an upward trend in the branching ratio (+0.0145 per year).
Table 3: Hawkes Process Parameters and Stationarity Tests
Parameter Estimate 95% CI
Baseline Intensity ($\mu$) 0.0159
Excitation Parameter ($\alpha$) 0.0223
Decay Rate ($\beta$) 0.0224
Branching Ratio (BR = $\alpha/\beta$) 0.9978 [0.050, 0.660]
Excludes Critical Value (BR=1.0)? Yes
Sample Characteristics:
Number of Surge Events 126
   Sell-Side Surges 116 (92%) (panic selling)
   Buy-Side Surges 10 (8%) (FOMO buying)
Observation Period (days) 1,259
Empirical Rate ($N/T$) 0.100
Stationarity and Equilibrium Tests:
Hurst Exponent (H) 0.817 (power-law decay)
ADF Test (p-value) 0.998 (non-stationary)
KPSS Test (p-value) 0.010 (non-stationary)
Yearly BR Trend (slope) +0.0145/year (trending toward criticality)
@@
High Herding Days 126 (10.0%)
Normal Days 1,133 (90.0%)
Note: Bootstrap confidence interval (1,000 iterations) excludes unity, confirming the process is statistically subcritical. Stationarity tests (ADF, KPSS) both reject stationarity, indicating non-equilibrium transient dynamics. The discrepancy between theoretical steady-state intensity ($\bar{\lambda}=\mu/(1-n)\approx 7.45$) and empirical rate (0.100) reflects transient regime behavior. H = 0.817 >> 0.55 indicates power-law decay characteristic of near-critical dynamics.
@@
The directional composition is strikingly asymmetric: of 126 surge events, 116 (92%) are sell-side and only 10 (8%) are buy-side. Retail herding in our sample is predominantly a panic-selling phenomenon. The explosive dynamics we observe (unconstrained BR = 1.09; Appendix A.1) reflect market-wide liquidation cascades rather than FOMO-driven buying frenzies.
Note that the theoretical steady-state intensity ($\bar{\lambda}=\mu/(1-n)\approx 7.45$ events/day) far exceeds the observed rate (0.100 events/day). This discrepancy reflects transient dynamics; we explore this in detail in Section 5.3.
Figure 2 shows the regime-conditional impulse response functions. During normal periods (blue lines), Foreign and Institutional investors exhibit the expected positive, persistent impact. But during high herding periods (red lines), their impact collapses and turns negative.
Figure 2: The Breakdown of Signal During Retail Herding. Panel A (Hawkes Intensity): Classifies market days into “Normal” (90%) and “High Herding” (10%) regimes based on the self-excitation intensity of individual investors. Panel B (Conditional Kernels): Shows the cumulative price impact of Institutional flows in both regimes. In “Normal” conditions (blue line), impact is positive and persistent (+0.0047). In “High Herding” conditions (red line), impact collapses and becomes deeply negative (–0.0447), indicating regime-dependent changes in institutional impact.
@@
Panel C: Size Quintile Breakdown (Institutional Impact by Market Cap)
[…]
@@
Note: Panel A reports the mean of stock-level estimates (equal weight per stock). Panel B reports estimates from pooled regressions (weight proportional to observations/liquidity). Panel C reveals that impact deterioration during herding is size-dependent: Q1-Q2 (40% of stocks) show persistently negative impact that worsens during herding; Q3 (20%) shows improvement (ratio << 1); Q4-Q5 (40%) maintain positive impact that amplifies during herding.
The key implication is that herding amplifies a pre-existing size gradient rather than creating a new one. Small-cap stocks, where institutional flows already carry negative impact—likely reflecting adverse selection or thin liquidity—deteriorate further during herding as retail noise intensifies. Large-cap stocks, with deeper liquidity and greater transparency, not only maintain but strengthen positive institutional impact during the same episodes. The market does not uniformly become opaque; rather, the cross-sectional dispersion in efficiency widens.